# Where MLLMs Attend and What They Rely On: Explaining Autoregressive Token Generation

### v2

@@ 4.5 Ablation Study
5 Conclusion and Limitation
In this paper, we present Eagle, a black-box attribution framework for autoregressive MLLMs. By unifying sufficiency and indispensability in a submodular-inspired objective, Eagle faithfully explains token generation, revealing both where models attend and what they rely on. Experiments across diverse models and datasets show clear gains in faithfulness, localization, and hallucination diagnosis. Moreover, by identifying and removing minimal interfering regions, Eagle also mitigates hallucinations, serving as both an interpretability and correction-oriented tool.
Limitations.
Despite its effectiveness, our work has two main limitations. First, the iterative subset selection and greedy search limit scalability compared to lightweight visualization methods. Second, the framework focuses on hallucination explanation and partial mitigation, leaving proactive prevention unexplored. Future work will explore faster search strategies and explanation-guided debiasing for training MLLMs.
Appendix A LLM Usage
During the preparation of this manuscript, large language models (LLMs) were employed in a limited and auxiliary capacity. Specifically, their usage was restricted to the following three aspects: (1) checking grammar and expression at the sentence level, thereby providing local linguistic refinement; (2) performing global polishing after the draft was completed, ensuring that the overall exposition conforms to idiomatic English usage.
At no stage were LLMs used for generating research ideas, developing arguments, or modifying the substantive content of this work. Their sole role was to assist in enhancing the clarity and effectiveness of communication.
Appendix B Eagle Algorithm
@@ 5 $\alpha\leftarrow\arg\max_{\alpha\in S_{d}}\,\mathcal{F}\!\left(\pi\cup\{\alpha\}\right)$;
Appendix C Additional Experimental Details
@@
Appendix D Additional Qualitative Results
@@
D.1 Sentence-level Explanations on MS COCO and MMVP
@@
D.2 Object-level Explanations on MS COCO
Beyond sentence-level results, we further evaluate our method at the object level with ground-truth bounding boxes. Fig. 12, Fig. 13, and Fig. 14 illustrate that our method produces sparse yet highly accurate localization of queried objects such as “boat,” “keyboard,” or “truck.” By contrast, IGOS++ frequently covers overly broad regions, while LLaVA-CAM and TAM often fail to precisely localize objects. These comparisons highlight the advantage of our method in generating interpretable, object-centric attributions.

### v4

@@ 4.5 Ablation Study
5 Conclusion
In this paper, we present Eagle, a black-box attribution framework for autoregressive MLLMs. By unifying sufficiency and indispensability in a submodular-inspired objective, Eagle faithfully explains token generation, revealing both where models attend and what they rely on. Experiments across diverse models and datasets show clear gains in faithfulness, localization, and hallucination diagnosis. Moreover, Eagle pinpoints the minimal visual factors that give rise to hallucinations, providing an effective, interpretation-focused account of their causes.
Supplementary Material
6 Eagle Algorithm
@@ 5  $\alpha\leftarrow\arg\max_{\alpha\in S_{d}}\,\mathcal{F}\!\left(\pi\cup\{\alpha\}\right)$;
7 Additional Experimental Details
@@
8 Limitations and Future Works.
Despite its effectiveness, our work has two main limitations. First, the iterative subset selection and greedy search limit scalability compared with lightweight visualization methods. Nevertheless, our approach provides a promising interpretable pathway and clarifies a potential upper bound of attribution for MLLMs; in future work, we will design more efficient attribution algorithms. Second, the framework focuses on hallucination explanation and partial mitigation, leaving proactive prevention unexplored. In future work, we will leverage explanations to automatically detect and mitigate hallucinations and, once failure modes are identified, develop data-/parameter-efficient methods for minimal-cost model repair.
9 Additional Qualitative Results
@@
9.1 Sentence-level Explanations on MS COCO and MMVP
@@
9.2 Word-level Explanations on MS COCO
Beyond sentence-level results, we further evaluate our method at the word-level with ground-truth bounding boxes. Fig. 12, Fig. 13, and Fig. 14 illustrate that our method produces sparse yet highly accurate localization of queried objects such as “boat,” “keyboard,” or “truck.” By contrast, IGOS++ frequently covers overly broad regions, while LLaVA-CAM and TAM often fail to precisely localize objects. These comparisons highlight the advantage of our method in generating interpretable, object-centric attributions.
9.3 Additional Hallucination Attribution Visualizations